OpenStudy (anonymous):

What number must you add to complete the square? x^2+4x=3

OpenStudy (anonymous):

add 4

OpenStudy (uri):

x^2+4x-3=0

OpenStudy (anonymous):

So it would be -3?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

I think so.

OpenStudy (lncognlto):

Well you could use the quadratic formula, \[-b \pm \sqrt{b ^{2}-4ac}\div2a\], where the equation is in the form ax^2 + bx + c =0, but completing the square is quicker in this case.

OpenStudy (anonymous):

I keep getting a Decimal.

OpenStudy (uri):

http://blogs.discovermagazine.com/loom/files/2008/07/quadratic.jpg Quad formula a:1 b:4 c:-3

OpenStudy (lncognlto):

4/2 = 2 2^2 = 4

OpenStudy (anonymous):

So 2^2 = 4 and the answer is 4,Correct?

OpenStudy (anonymous):

yeah

OpenStudy (lncognlto):

The answer to the question is not 4, but 4 is the number added to the equation. Here are step by step instructions for completing squares: http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (lncognlto):

There should be two answers to the equation.

OpenStudy (anonymous):

there is

OpenStudy (uri):